OpenStudy (anonymous):

Quick question before quiz: How do i find pi/12?

OpenStudy (anonymous):

Hello again Nnesha

Nnesha (nnesha):

divide that into two angle what two radian on the unit circle add up to pi/12

Nnesha (nnesha):

heya! :=))

OpenStudy (anonymous):

ummm, I actually have no Idea

Nnesha (nnesha):

btw it's just pi/12 or sin(pi/12) is there any ratio with pi/12 ?

Nnesha (nnesha):

i guess we have to use addition formulas

OpenStudy (anonymous):

O_O

Nnesha (nnesha):

hmm what ???

OpenStudy (anonymous):

would a ratio be pi/6?

Nnesha (nnesha):

pi/6+ what = pi/12

Nnesha (nnesha):

btw you can subtract or add doesn't matter

Nnesha (nnesha):

\[\frac{ \pi }{ 12 } = \frac{\pi}{4} - \frac{\pi}{6}\]

Nnesha (nnesha):

right pi/4 - pi/6=pi/12 and both angle pi/4 and pi/6 are on the unit circle

OpenStudy (anonymous):

how does that happen? pi/4+pi/6 get me pi/12

Nnesha (nnesha):

what's the common denominator \[\frac{ \pi }{ 4 } +\frac{\pi}{6}\] ??

OpenStudy (anonymous):

6

OpenStudy (anonymous):

or 12 ^_^

Nnesha (nnesha):

hmm it's not 6 if the denominators aren't the same then we should multiply them \[\huge\rm \frac{ a }{ \color{ReD}{b} }+\frac{c}{\color{blue}{d}}\] \[\rm \frac{ ??+?? }{ bd }\] when we find common denominator we should multiply -numerator of first fraction by the denominator of 2nd fraction -multiply numerator fo 2nd fraction by the denominator of first fraction \[\rm \frac{ a\color{blue}{d}+c\color{red}{b} }{ bd }\]

OpenStudy (anonymous):

so with that formula, we get pi/24

Nnesha (nnesha):

hmm \[\large\rm \frac{ \pi }{ \color{Red}{4} } +\frac{\pi}{\color{blue}{6}}\] \[\rm \frac{ \color{blue}{6}\pi+\color{red}{4}\pi }{ 24 }\] multiply the numerator of first fraction by the denominator of 2nd fraction and multiply multiply the numerator of 2nd fraction by the denominator of first fraction in other words cross multiply there is another way to find common denominator